Oxford Materials IT Committee Meeting


A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 27th October 2015 in the PR Meeting Room


  • 1 Apologies for absence and Membership of the Committee
  • 2 Matters outstanding from the previous meeting
  • 3 Matters for discussion (attached)
  • 4 Head of IT report (attached)
  • 5 Finances
  • 6 AOB
  • 7 Date and time of next meeting: Tuesday 2nd February 2016 at 11am in PR Committee Room


See minutes below.


IT services projects: RDM, TONE, Chorus, ISP, Innovation seedfund
Departmental projects: Photocopiers, Green Impact, iLab
Information from other Committees: MPLS ICT Panel, Divisional Software Licensing Group


The report below summarises activities since the last meeting of ITC, as categorised by topic. The report is intended for record rather than for discussion at ITC meeting although queries may arise.

Personnel: Khalid Schofield’s grade 5 position was mothballed and a new position created which was graded as 6 (rather than the desired 7). The position was advertised but the number of applicants was low and the position was ultimately unfilled despite being offered. The position will be re-advertised.

Admin Network Replaced desktops for Philippa Moss, Ashley Brown, Marion Beckett and Patrick Grant. Numerous rearrangements associated with refurbishment of Head of Department office. Sarolta Mohaine Palfi joined as PA to Harish Bhaskaran and Chris Grovenor. Several temporary storesmen were added and removed. Oracle JRE updated for HRIS.

Teaching LaboratoryComputer provision was reviewed over the summer but not upgrades were required since all systems are less than 2 years old.

Lecture theatres / meeting rooms One visualizer was repaired (replacement psu). A small meeting area was implemented in the Head of Department’s office. There are still plans for Skype functionality in ETBCR and BRCR plus electronic display board for visitors both in HR reception and also at Begbroke.

Digital Print-room No significant changes or issues.

Library The self-issue system still needs upgrading. New wireless access point was installed. Usage of the “hot desk” pc requested by graduates seems negligible.

MML Rothery2 was decomissioned. Zener was merged with Landau. Ironman continues to have issues. Aircon failed and was repaired twice over the summer – systems were down for 2 weeks...

FIM New atom probe is being installed. Plans for replacement of image acquisition computer on old 3DAP.

EM WinSCP+putty was rolled out across the department to replace Secure Shell client.

MFFP YSeveral failed disk replaced in the TJM storage array. A RAID crash was recovered after two weeks which caused review of the lack of backup and implementation of duplication of the most important data (although overall the arrays are too large to back up). Monthly patching and rebooting of 13 systems and storage arrays. Office space reorganisation and associated rearrangement of computers.

OMCS Nicola Flanagan and Julie Page started.

Rex Richards / Peter Bruce Group Salto door entry system still needs fixing to work with university cards rather than separate cards. Steady stream of new people and equipment.

Website CMSmadesimple upgraded to 1.13 for all sites. Daily/weekly editing of department website and annual refresh for new academic year. MFFP site needs migrating to drupal 7. New webserver needs implementing and all sites migrating to https. CMSmadesimple sites need upgrading to version 2 within a year.

Publications 9 sets of business cards. Printed handbooks for new year (still waiting for new version of Finals handbook and final version of Graduate handbook). Weekly editing of Departmental Newsletter. Signage and posterboards for opendays and four conferences/workshops (MFFP,QIP,MML,NG).

Network Netman was decommissioned. Twenty-seven gigabit network switches were purchased as budgeted but not rolled-out yet due to lack of manpower. The 5 wireless access points in HR were replaced with unifi access points and the old access points reallocated elsewhere.

Security Ashley Brown has taken over management of the Kemet door entry. A security door entry camera was implemented at Rex Richards with recording. Oxcert have released 61 security bulletins. The department webserver security was reviewed by an external company (paid centrally) and generally we passed but with recommendation that most websites should migrate to https. There have been a handful of browser-hijacks and virus infections. The BR door entry network control unit was upgraded to fix intermittent drop-out.

WDE Still only six user’ devices encrypted – need to expand uptake targeting senior staff. Investigated FileVault for Mac but ultimately re-encrypted Alana’s laptop with WDE.

DisposalWEEE disposal happened and two van loads of equipment were disposed, including clearing items from Khalid Schofield’s and Richard Cripps’s offices.

Software Old management server oums-nomad was decommissioned. Windows 10 was tested prior to official release. We are able to deploy Windows 10 but still prefer to deploy Windows 7. The few existing Windows8 systems are being recommended to upgrade to Windows10. Office 2016 was released.

Disaster Recovery and Backup Several disk fails have occurred and data has been recovered or restored from backups.

Documentation updates Documentation on the private Mediawiki continues to expand, but time for documentation has been limited due to being short staffed.

New people & accounts May 7; June 11; July 8; August 10; September 40; October 39;

New Device registrations May 20; June 8; July 17; August 6; September 24; October 20;


A meeting of the Department of Materials IT Committee was held at 11am on Tuesday 27th October 2015 in the PR Meeting Room. The following people attended: APD, PJW, RS, RT, SPF, PAJB, Flynn Castles, Andy Naylor, Lewys Jones, Qifeng Yang

1. Apologies for Absence and Membership of the Committee

Membership of all departmental committees was reviewed over the summer. The committee welcomed Steven Fitzgerald and Paul Bagot as academic representatives and thanked Simon Benjamin and Jonathan Yates for their contributions over many years.

Qifeng Yang attended as JCCG rep on behalf of Karwei So.

Chris Akinola sent his apologies.

2. Matters arising from the minutes

(Items marked with * were not for discussion at the meeting)

a) *Draft Information Security Policy (Minute 2b)
Again there has been little opportunity to progress towards formalising the departmental Information Security Policy for approval by the ITC and subsequently DC. However PJW noted that the ISP documentation and toolkit has evolved considerably over the last few months and PJW is keen to revisit implementing a departmental IS policy.

Action: PJW

b) Finances (Minute 5)
Paul journalled costs for stock computers from August 2014 to July 2015. A couple of grant codes had ended and were journalled elsewhere.

c) Office2013 for Admin (Minute 6a)
APD asked whether Office2013 could be rolled out more widely. PJW to investigate compatibility with central university systems such as macros in R12 reports.

Action: PJW

3 Matters for Discussion

a) Verbal update on various projects

PJW updated the committee verbally upon the status of the following projects:

  • RDM http://researchdata.ox.ac.uk
    ESPRC guidelines on publishing all research data came into force on 15th May 2015. The university policy only requires publication of data supporting the publication itself (graphs, images, databases etc.) rather than all the underlying raw and analysed data for a research project. Patrick has taken over as Open Access coordinator for the department and will be reviewing procedures.
  • TONE – replacement backbone network. http://projects.it.ox.ac.uk/tone/
    Core backbone fibre is in place. Testing the core switching equipment encountered problems during load testing so the pilot has been delayed until January 2016.
    Ultimately there will be no effect upon departmental users since the Department feed will remain at 1Gb backbone connection. Several buildings have had replacement fibres feeds installed (12/13 PR, 14PR, 21BR, HB, ETB) however our main feed in HR cannot be replaced due to asbestos. Other buildings RR and BB are shared buildings and are still awaiting new fibre feeds.
  • Chorus – Chorus – the new telephone system http://projects.it.ox.ac.uk/icp/. Also see official Chorus newsletters http://projects.it.ox.ac.uk/icp/newsletter
    This is a big project involving revamping all phones within the Department. The new phone system is number-per-person and a software based system which can redirect phone calls to any number of devices (phone, computer, smartphone etc.) and will require a change in user interaction. In addition the project will force significant upgrade and expansion of the Department network as the phone network is converged with the data network. 28 new switches have been purchased are being deployed. The June pilot project uncovered some issues and the pilot will be re-run.
  • Innovation Seed Fund – Browse and contribute at https://oxfordideas.wazoku.com/. The third round will close at end November. [Topics: undergrad admissions and access, facilitating research, international collaboration, ideas for reducing carbon emissions]. PJW is keeping a watching brief and will alert members of department as considered necessary.
  • C3E – https://sharepoint.nexus.ox.ac.uk/sites/itservices/c3e
    The university has officially agreed to plan to migrate the in-house Nexus Exchange service to a cloud-based Microsoft Office365 service. The Nexus Sharepoint service will remain in-house. This migration will take several years to complete.
  • Lecture capture project – http://help.it/ox.ac.uk/replay
    This department does not have the manpower to actively engage with this new initiative on recording of lectures for students.
  • Java Runtime Environment Upgrade
    HRIS has been upgraded to v20 on 27th October and now uses JRE 1.8. However Financials and Oracle Student System still requires JRE 1.6 although these are also due for upgrade in the coming months. JRE 1.6 support ends in December 2015.
  • Information Security Improvement Programme (ISIP)
    This department underwent an external webserver security audit as part of the ISIP project. PJW will move forward with some recommendations while commissioning a new department webserver. This department still needs to review and adopt an Information Security Policy.(Minute 2a) A new ISP website is due to be launched before 2016.

Departmental projects:

  • Photocopiers – The department photocopiers were replaced in June with new colour photocopiers that can all act as printers and scanners. The department website has been updated. PJW will advertise via newsletter and monitor usage monthly.
  • iLab – The system has been running for one year. APD has been talking with iLab about further refinement of reporting and the need for collecting actual usage data. [APD requested PJW keep reporting upon iLab at ITC]
  • Green Impact – the Green Impact team submitted the workbook on the 17th April and were awarded Silver for the Department and Gold for the Labs- well done!. PJW is still actively involved and is keen to renew interest and keep actively improving recycling, power saving, low carbon etc.

Information from other committees:

  • MPLS ICT panel
    the Michaelmas Term meeting yesterday received a status report from IT Services (mostly covered above). Officially the MPLS ICT panel should have representation from each department of an IT manager and an academic, however currently only PJW attends on behalf of Materials.
  • Divisional Software Licensing Group
    Matlab site license was renewed paid centrally by IT Services but with an expectation of divisional contributions. Windows 10 and Office2016 are available.

4. Head of IT Report
A written report was circulated with the agenda.

5. Finances
As budgeted, £20k was spent in July and August purchasing 28 network switches in order to upgrade the department network to entirely gigabit. PJW is up to date in terms of journaling costs for stock desktop computers since 1st August.

6. AOB

No other business.

7. Date of Next Meeting
The next meeting is Tuesday 2nd February 2016 at 11am in PR Meeting Room.

The meeting closed at 12.05pm.